Compound Amount


Abraham ace invested $200,000 at 14% interest compounded semi-annually for 8 years. What is the amount of compound interest Abraham earned.


1
Expert's answer
2021-03-23T02:37:07-0400

"Principal = 200,000"

"Amount=p(1+R\/n)"nt

P-rincipal

R-interest rate.

n-number of times interest applied per time period.

nt-number of time periods elapsed.

"A=200,000(1+0.14\/2)"2*8

"=200,000(2.952163749)"

$"590,432.7497"

Compound interest

"=590,432.7497-200,000"

"=$390,432.7497"